OpenStudy (anonymous):

How do you solve 12n + 2 - 3n = 10 - n +2

OpenStudy (jadeishere):

add the like terms and isolate th variable (n) 12n +2 - 3n = 10 - n + 2 9n + 2 = 10 -n + 2 10n = 10 n = 1

OpenStudy (anonymous):

I do not understand I'm at 9n + 2 = 10 - n + 2

OpenStudy (jadeishere):

oh sorry /)_- 9n + 2 = 10 - n + 2 so take the - n from the right side and do the opposite equation on both sides (+n) so (+n) +9n + 2 = 10 - n + (n) + 2 so 10n + 2 = 10 + 2 then move over the 2 to the right side 10n +2 - (2) = 12 - (2) 10n = 10

OpenStudy (anonymous):

YOU'RE NOT MAKING SENSE BYE!!!

OpenStudy (jadeishere):

idk how to explain it.. What's not making sense?

OpenStudy (texaschic101):

do you understand what like terms are ?

OpenStudy (texaschic101):

12n + 2 - 3n = 10 - n + 2 --- combine like terms 9n + 2 = 12 - n -- add n to both sides 9n + n + 2 = 12 - n + n -- combine like terms 10n + 2 = 12 -- subtract 2 from both sides 10n + 2 - 2 = 12 - 2 -- combine like terms 10n = 10 -- divide both sides by 10 10n/10 = 10/10 n = 1

OpenStudy (texaschic101):

I do not know how to explain it any easier....

OpenStudy (jadeishere):

Neither do I!!